Am ordinary(fair) die is a cube with the numbers 1 thru 6 on the sides ( representEd by painted spots). Imagine that such a die is rolled twice in succession and that the face values of the two rolls are added together. This summer is recorded as the outcome of a single trial of a random experiment.
Compute the probability of each of the following events:
Event A: the sum is greater than 7
Event B: themail sum is an odd number
Write your answers as exact fractions.

Answer by emathwiz(15)

For this kind of problem the probability is the number of outcomes that work over the total number of outcomes. The total number of outcomes for two rolls is
6+5+4+3+2+1=21 This works because:
(first roll, second roll)
(1,1) (1,2) (1,3) (1,4) (1,5) (1,6)= 6
(2,2) (2,3) (2,4) (2,5) (2,6)=5
(3,3) (3,4) (3,5) (3,6)=4
(4,4) (4,5) (4,6)=3
(5,5) (5,6)=2
(6,6)=1


Looking at all these possibilities from here we just count the ones that work.
9 sums are greater than 7 so the probability for Event A is 9/21 or 3/7
For Event B there are once again 9 odd sums so the probability is 9/21 or 3/7.
Keep in mind this is if order doesn't matter! (That's why I don't include 6,3 for example since I wrote 3,6 already)